### v1

@@ 5 Model dependence
Before analyzing the experimental data, we investigated the dependence of neural network predictions on the choice of high-energy hadronic interaction model. To this end, we trained two neural networks: one using events simulated with QGSJET-II-04 and EPOS-LHC as high-energy hadronic interaction models and the other using all available models (including Sibyll). For this study, we did not perform fine-tuning of neural networks on experimental data.
@@
We also examined the model dependence of neural network predictions for photon-induced air showers. Figure 7 illustrates the histograms of the neural network predictions for Monte Carlo events simulated using QGSJET-II-04 and EPOS-LHC. The histograms exhibit a discrepancy for proton-like events ($\xi<0.1$) but are in good agreement for $\xi>0.5$. This observation shows that the evolution of photon-induced air showers exhibits some dependence on the choice of hadronic interaction model. However, for non-proton-like events ($\xi>0.5$), which are of primary interest for this study, the difference is negligible. This justifies our decision to use only one hadronic interaction model for simulating photon-induced air showers.
@@ 6 Results
As an additional validation measure, we examined the neural network’s predictions on events correlated with lightning strikes registered near TA SD, Figure 8(a). The corresponding histogram exhibits a peak near $\xi=1$, indicating that some of the events are indeed lightning-induced. We have also identified Terrestrial Gamma-Ray Flashes and lightning-correlated events previously reported by TA as photon-like with high confidence [kieu2024first, Abbasi:2024vlj, TelescopeArrayProject:2017elp]. This confirms that the neural network correctly learned to identify photon-like events since lightning-induced events are essentially electromagnetic cascades, which are similar to photon-induced air showers [kieu2024first].
@@
Figure 9: Upper limits on diffuse flux of UHE photons obtained in this study (TA 14 yr) compared with the previous TA result [TelescopeArray:2018rbt] and limits obtained by AGASA [Shinozaki:2002ve], Yakutsk [Glushkov:2009tn], and Pierre Auger hybrid [PierreAuger:2024ayl] and SD [abreu2023search]. For comparison we show the expected flux of cosmogenic UHE photons for several UHECR flux models: pure protons (red band [Hooper:2010ze] and hatched band [Gelmini:2022evy]), pure iron (blue band) [Hooper:2010ze], in both cases the bands denote the theoretical uncertainties of the models, and TA best-fit model [Bergman:2021djm, Kuznetsov:2025ehl]. The predictions for SHDM decaying via hadronic and leptonic channels are also shown [Kachelriess:2018rty].
Energy, log10(E/1eV) Quality cuts $\xi$-cut $\xi$-cut value Effective exposure,
$10^{3}~\text{km}^{2}\cdot\text{sr}\cdot\text{yr}$
@@
Energy, log10(E/eV) Effective exposure, $10^{3}~\text{km}^{2}\cdot\text{sr}\cdot\text{yr}$ Proton suppression level Expected photon candidates (from proton bg.) Observed photon candidates Photon flux upper limit, $(\text{km}^{2}\cdot\text{sr}\cdot\text{yr})^{-1}$ Photon fraction upper limit
@@
Table 2: Resulting metrics and limits on the photon flux. Effective exposure is evaluated after imposing both the event quality cuts and the $\xi$ cut.
To derive upper limits on the photon flux, we use Eq. 2.1. We calculate the effective TA SD exposure taking into account TA SD geometry, event selection and trigger efficiency, exclusion of events below the classification threshold, and the use of burn sample for fine-tuning the neural network. The efficiency of quality and $\xi$ cuts along with the effective exposure is shown in Table 1.
The resulting limits on photon flux are presented in Table 2. We observed no significant excess of photon candidates compared to the expected proton background. The upper limits on the photon fraction were calculated as a ratio of the corresponding photon flux limit to the integral flux of the TA SD spectrum [TelescopeArray:2023bdy]. In Fig. 9 we compare the present limits with the limits of AGASA [Shinozaki:2002ve], Yakutsk [Glushkov:2009tn] and Pierre Auger hybrid  [PierreAuger:2024ayl] and SD observations [abreu2023search]. For comparison we depict the previous TA limits obtained using boosted decision trees for 9 years of SD operation time [TelescopeArray:2018rbt]. In the same figure we also show theoretical predictions for the UHE photon flux. The predictions for cosmogenic photons (GZK photons) depend on the mass composition of initial UHECR flux: pure protons [Hooper:2010ze, Gelmini:2022evy] and pure iron nuclei [Hooper:2010ze] are shown as extreme cases. The uncertainties of the predicted photon fluxes due to uncertainties of the initial UHECR flux and its propagation are shown as bands. We also show the prediction of the cosmogenic UHE photon flux [Kuznetsov:2025ehl] for the UHECR flux model derived as a best fit for the TA spectrum and $X_{\rm max}$ data [Bergman:2021djm]. The expectations for UHE photon fluxes from the decay of super heavy dark matter (SHDM) particles $X$ are shown for two benchmark models: $X\rightarrow q\bar{q}$ and $X\rightarrow\nu\bar{\nu}$ [Kachelriess:2018rty]. In both cases the $X$ mass, $M_{X}$, is $10^{12}$ eV and the its lifetime, $\tau_{X}$, is normalized to the Pierre Auger hybrid limits [PierreAuger:2024ayl], which are the most stringent constraints for these scenarios. Specifically, we use $\tau_{X}=7.1\times 10^{22}$ yr for the $X\rightarrow q\bar{q}$ model and $\tau_{X}=1.7\times 10^{22}$ yr for the $X\rightarrow\nu\bar{\nu}$ model.
@@ 7 Conclusions
We found no significant excess of photon candidates above the expected hadronic background. Our limits approach the upper edge of the GZK photon predictions [Hooper:2010ze, Gelmini:2022evy]; however, probing the GZK photon models [Kuznetsov:2025ehl] will require an improvement of several times in sensitivity. Notably, in the search for signals from SHDM decay, our limits are competitive with those of Auger, despite the fact that TA does not observe observe the Galactic center region, which is abundant for the dark matter. Looking ahead, we hope to support or constrain such models in future with increased TA SD exposure time or improved machine learning-based event selection criteria.
@@ Appendix A Reconstruction procedure
The reconstruction procedure employs a joint fit of the event geometry and lateral distribution function [TelescopeArray:2013yze]. The fit incorporates seven free parameters:
@@
where $r$ is the distance from the station to the shower core (i.e. the shortest distance between a point and a line), $t_{\text{plane}}(\vec{r})$ is the arrival time of the shower’s plane front to the station located at $\vec{r}$, and $LDF(r)$ is the empirical lateral distribution profile introduced in the AGASA experiment [teshima1986properties] and fine-tuned to better fit TA data [abu2013cosmic]:
@@
The energy of the primary particle is estimated as a function of $S_{800},~\theta$, and $\phi$ using a lookup table obtained from the photon MC simulations [TelescopeArray:2018rbt].
Below is the list of 14 reconstructed parameters [TelescopeArray:2018rbt] used as input data for the neural network:
@@
    Area-over-peak of the signal at 1200 m and the corresponding slope parameter [abraham2008upper];
@@
    $S_{b}\equiv\sum_{i}\left(S_{i}\times\left(\frac{r_{i}}{r_{0}}\right)^{b}\right)$ for two values of $b$: 2.5 and 4.0. Here $S_{i}$ and $r_{i}$ are the integral signal of $i$-th station and its distance to the shower core, respectively, and $r_{0}=1200~m$ [Ros:2011zg];
@@
    Sum of integral signals from all of the triggered stations.
@@
    Asymmetry of the signal between the upper and lower layers of stations [abbasi2019mass];
@@
    Total number of peaks within all stations [abbasi2019mass];
@@
    Minimal distance between the reconstructed air shower core and SD array edge.

### v2

@@ 5 Model dependence
QGSJET-II-04, EPOS-LHC, and Sibyll differently approximate high-energy hadron interactions, which, in particular, results in different muons yield. Since higher muon yield would make easier to distinguish proton-induced and photon-induced events, we investigated the dependence of neural network predictions on the choice of high-energy hadronic interaction model. To this end, we trained two neural networks: one using events simulated with QGSJET-II-04 and EPOS-LHC as high-energy hadronic interaction models and the other using all available models (including Sibyll). For this study, we did not perform fine-tuning of neural networks on experimental data.
@@
We also examined the model dependence of neural network predictions for photon-induced air showers. Figure 7 illustrates the histograms of the neural network predictions, trained on QGSJET-II-04 MC only, for the events simulated using QGSJET-II-04 and EPOS-LHC. The histograms exhibit a discrepancy for proton-like events ($\xi<0.1$) but are in good agreement for $\xi>0.5$. This observation shows that the evolution of photon-induced air showers exhibits some dependence on the choice of hadronic interaction model. However, for non-proton-like events ($\xi>0.5$), which are of primary interest for this study, the difference is negligible. We also tested that in the high photon-likeness region, $\xi>0.9$, photon exposure difference is below 7% for all energy ranges. This justifies our decision to use only one hadronic interaction model for simulating photon-induced air showers.
@@ 6 Results
As an additional validation measure, we examined the neural network’s predictions on events correlated with lightning strikes registered near TA SD, Figure 8(a). The corresponding histogram exhibits a peak near $\xi=1$, indicating that some of the events are indeed lightning-induced. We have also identified Terrestrial Gamma-Ray Flashes and lightning-correlated events previously reported by TA as photon-like with high confidence [60, 8, 5]. This confirms that the neural network correctly learned to identify photon-like events since lightning-induced events are essentially electromagnetic cascades, which are similar to photon-induced air showers [60].
@@
Figure 9: Upper limits on diffuse flux of UHE photons obtained in this study (TA 14 yr) compared with the previous TA result [6] and limits obtained by AGASA [77], Yakutsk [43], and Pierre Auger hybrid [11] and SD [15]. For comparison we show the expected flux of cosmogenic UHE photons for several UHECR flux models: pure protons (red band [47] and hatched band [41]), pure iron (blue band) [47], in both cases the bands denote the theoretical uncertainties of the models, and TA best-fit model [27, 67]. The predictions for SHDM decaying via hadronic and leptonic channels are also shown [51].
Energy, log10(E/1eV) Quality cuts Photon selection efficiency, $S(\xi)$ $\xi$-cut value Effective exposure, $10^{3}~\text{km}^{2}\cdot\text{sr}\cdot\text{yr}$
@@
Energy, log10(E/eV) Proton suppression level Expected proton bg. Observed photon candidates $\Phi_{\gamma}$, $(\text{km}^{2}\cdot\text{sr}\cdot\text{yr})^{-1}$ Photon fraction upper limit
@@
Table 2: Resulting metrics and limits on the photon flux $\Phi_{\gamma}$. Effective exposure is evaluated after imposing both the event quality cuts and the $\xi$ cut. Proton suppression level is calculated by directly counting the number of proton-induced events below and above the $\xi$ cut.
To derive upper limits on the photon flux $\Phi_{\gamma}$, we use Eq. 2.1. We calculate the effective TA SD exposure taking into account TA SD geometry, event selection and trigger efficiency, exclusion of events below the classification threshold, and the use of burn sample for fine-tuning the neural network. The efficiency of quality and $\xi$ cuts along with the effective exposure is shown in Table 1.
The resulting limits on photon flux $\Phi_{\gamma}$ are presented in Table 2. We observed no significant excess of photon candidates compared to the expected proton background. The upper limits on the photon fraction were calculated as a ratio of the corresponding photon flux limit to the integral flux of the TA SD spectrum [61]. The energy, zenith angle and Linsley front curvature of the observed photon candidates are reported in Appendix C.
In Fig. 9 we compare the present limits with the limits of AGASA [77], Yakutsk [43] and Pierre Auger hybrid  [11] and SD observations [15]. For comparison we depict the previous TA limits obtained using boosted decision trees for 9 years of SD operation time [6]. In the same figure we also show theoretical predictions for the UHE photon flux. The predictions for cosmogenic photons (GZK photons) depend on the mass composition of initial UHECR flux: pure protons [47, 41] and pure iron nuclei [47] are shown as extreme cases. The uncertainties of the predicted photon fluxes due to uncertainties of the initial UHECR flux and its propagation are shown as bands. We also show the prediction of the cosmogenic UHE photon flux [67] for the UHECR flux model derived as a best fit for the TA spectrum and $X_{\rm max}$ data [27]. The expectations for UHE photon fluxes from the decay of super heavy dark matter (SHDM) particles $X$ are shown for two benchmark models: $X\rightarrow q\bar{q}$ and $X\rightarrow\nu\bar{\nu}$ [51]. In both cases the $X$ mass, $M_{X}$, is $10^{12}$ eV and the its lifetime, $\tau_{X}$, is normalized to the Pierre Auger hybrid limits [11], which are the most stringent constraints for these scenarios. Specifically, we use $\tau_{X}=7.1\times 10^{22}$ yr for the $X\rightarrow q\bar{q}$ model and $\tau_{X}=1.7\times 10^{22}$ yr for the $X\rightarrow\nu\bar{\nu}$ model.
@@ 7 Conclusions
We found no significant excess of photon candidates above the expected hadronic background. Our limits approach the upper edge of the GZK photon predictions [47, 41]; however, probing the GZK photon models [67] will require an improvement of several times in sensitivity. Notably, in the search for signals from SHDM decay, our limits are competitive with those of Auger, despite the fact that TA does not observe observe the Galactic center region, which is abundant for the dark matter. Looking ahead, we hope to support or constrain such models in future with increased TA SD exposure time or improved machine learning-based event selection criteria.
@@ Appendix A Reconstruction procedure
The reconstruction procedure employs a joint fit of the event geometry and lateral distribution function [17]. The fit incorporates seven free parameters:
@@
where $r$ is the distance from the station to the shower core (i.e. the shortest distance between a point and a line), $t_{\text{plane}}(\vec{r})$ is the arrival time of the shower’s plane front to the station located at $\vec{r}$, and $LDF(r)$ is the empirical lateral distribution profile introduced in the AGASA experiment [81] and fine-tuned to better fit TA data [19]:
@@
The energy of the primary particle is estimated as a function of $S_{800},~\theta$, and $\phi$ using a lookup table obtained from the photon MC simulations [6].
Below is the list of 14 reconstructed parameters [6] used as input data for the neural network:
@@
    Area-over-peak of the signal at 1200 m and the corresponding slope parameter [14];
@@
    $S_{b}\equiv\sum_{i}\left(S_{i}\times\left(\frac{r_{i}}{r_{0}}\right)^{b}\right)$ for two values of $b$: 2.5 and 4.0. Here $S_{i}$ and $r_{i}$ are the integral signal of $i$-th station and its distance to the shower core, respectively, and $r_{0}=1200~m$ [75], corresponding to TA SD stations spacing;
@@
    Sum of integral signals from all of the triggered stations;
@@
    Asymmetry of the signal between the upper and lower layers of stations [10];
@@
    Total number of peaks within all stations [10];
@@
    Minimal distance between the reconstructed air-shower core and SD array edge.
@@ Appendix B Results without fine-tuning
Appendix C Photon candidates
Below we present the main reconstructed parameters of the observed photon candidates. Note that their photon-likeness is very close to the $\xi$-cut value for $10^{19.5}$ eV limit.
Date (yyyy.mm.dd) Time
$\xi$-value $E_{\gamma}$, EeV $E_{p}$, EeV $\theta_{\gamma}$ $\phi_{\gamma}$ $a_{curv}$
2011.08.02 055738.387354 0.934 11.57 5.08 23.0° 208.7° 1.44
2011.10.02 235948.253047 0.939 37.29 13.57 15.1° 110.9° 0.75
2015.07.09 052937.805091 0.935 11.06 4.61 36.0° 11.6° 0.98
Table 3: Parameters of the observed photon candidates. $E_{\gamma}$ and $E_{p}$ are event energies reconstructed using photon-optimized and proton-optimized algorithms, $a_{curv}$ stand for Linsley front curvature. The reconstructed zenith and azimuth angles are given with respect to the standard TA SD coordinate system.